# Stacking Brick by Brick: Aligned Feature Isolation for Incremental Face Forgery Detection

### v2

@@ Abstract
The rapid advancement of face forgery techniques has introduced a growing variety of forgeries. Incremental Face Forgery Detection (IFFD), involving gradually adding new forgery data to fine-tune the previously trained model, has been introduced as a promising strategy to deal with evolving forgery methods. However, a naively trained IFFD model is prone to catastrophic forgetting when new forgeries are integrated, as treating all forgeries as a single “Fake” class in the Real/Fake classification can cause different forgery types overriding one another, thereby resulting in the forgetting of unique characteristics from earlier tasks and limiting the model’s effectiveness in learning forgery specificity and generality. In this paper, we propose to stack the latent feature distributions of previous and new tasks brick by brick, i.e., achieving aligned feature isolation. In this manner, we aim to preserve learned forgery information and accumulate new knowledge by minimizing distribution overriding, thereby mitigating catastrophic forgetting. To achieve this, we first introduce Sparse Uniform Replay (SUR) to obtain the representative subsets that could be treated as the uniformly sparse versions of the previous global distributions. We then propose a Latent-space Incremental Detector (LID) that leverages SUR data to isolate and align distributions. For evaluation, we construct a more advanced and comprehensive benchmark tailored for IFFD. The leading experimental results validate the superiority of our method.
@@ 3.4.1 Feature Isolation with Distribution Re-filling
$\mathcal{L}_{iso}=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{\exp(\mathbf{f}_{i}\cdot\mathbf{f}_{j}/\tau)}{\sum_{k=1}^{N}\mathbb{I}_{[y_{i}=y_{k}]}\exp(\mathbf{f}_{i}\cdot\mathbf{f}_{k}/\tau)}\right),$ (5)
where $\mathbf{f}_{i},\mathbf{f}_{j}$ are features from two different domains. $y_{i}$ is the domain label of $\mathbf{f}_{i}$, and it is allocated with a unique value to each real/fake and previous/new domain. $\mathbb{I}_{[y_{i}=y_{k}]}$ denotes an indicator function, which equals 1 if $y_{i}=y_{k}$, and 0 otherwise. Notably, $\mathbf{f}$ could be the feature of current training data if they are from the new task, and generated by SUR or re-filled data if they are from the previous tasks. Meanwhile, to encourage the learning of diverse real domains, the real data from different tasks is also assigned with different unique $y_{i}$.
@@ 3.4.2 Incremental Decision Alignment
$\mathbf{\theta}^{t+1}_{s+1}=\gamma\mathbf{\theta}^{t}_{s}+(1-\gamma)\frac{\mathbf{\theta}^{t+1}_{s}\|\mathbf{\theta}^{t}_{s}\|_{2}}{\|\mathbf{\theta}^{t+1}_{s}\|_{2}},$ (6)
where $s$ denotes the $s$-th step of optimization, and $\gamma$ denotes the learning rate. During training on the $(t+1)$-th task, the classifier $\mathcal{C}^{t+1}$ is optimized following Eq. 6 to be aligned with $\mathcal{C}^{t}$, while all previous classifiers are frozen to maintain the previous decision boundaries and their alignment.
@@ 3.5 Training and Inference
$y_{\text{infer}}=\sum_{i=1}^{t+1}\frac{\mathcal{C}^{i}(\mathbf{x})}{t+1}.$ (10)
Method Venue Task Protocol 1 Protocol 2
@@ 4.2 Comparisons with Existing Methods for Incremental Face Forgery Detection
Method FF++ DFDCP DFD CDF Avg.
LwF [22] 67.34 67.43 84.05 87.90 76.68
CoReD [19] 74.08 76.59 93.41 80.78 81.22
DFIL [29] 86.28 79.53 92.36 83.81 85.49
DMP [40] 91.61 84.86 91.81 91.67 89.99
Ours 90.89 89.33 93.97 94.34 92.13
Additionally, in Tab. 2, we copy the results after all tasks are incremented with P3 from their official papers [29, 40] to further compare the IFFD performance. Despite the notable distinction in experimental settings among these methods, our method still exhibits superior performance.

### v3

@@ Abstract
The rapid advancement of face forgery techniques has introduced a growing variety of forgeries. Incremental Face Forgery Detection (IFFD), involving gradually adding new forgery data to fine-tune the previously trained model, has been introduced as a promising strategy to deal with evolving forgery methods. However, a naively trained IFFD model is prone to catastrophic forgetting when new forgeries are integrated, as treating all forgeries as a single “Fake” class in the Real/Fake classification can cause different forgery types overriding one another, thereby resulting in the forgetting of unique characteristics from earlier tasks and limiting the model’s effectiveness in learning forgery specificity and generality. In this paper, we propose to stack the latent feature distributions of previous and new tasks brick by brick, i.e., achieving aligned feature isolation. In this manner, we aim to preserve learned forgery information and accumulate new knowledge by minimizing distribution overriding, thereby mitigating catastrophic forgetting. To achieve this, we first introduce Sparse Uniform Replay (SUR) to obtain the representative subsets that could be treated as the uniformly sparse versions of the previous global distributions. We then propose a Latent-space Incremental Detector (LID) that leverages SUR data to isolate and align distributions. For evaluation, we construct a more advanced and comprehensive benchmark tailored for IFFD. The leading experimental results validate the superiority of our method. Code is available at https://github.com/beautyremain/SUR-LID.
@@ 3.4.1 Feature Isolation with Distribution Re-filling
$\mathcal{L}_{iso}=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{\exp(\mathbf{f}_{i}\cdot\mathbf{f}_{j}/\tau)}{\sum_{k=1}^{N}\mathbb{I}_{[y_{i}\neq y_{k}]}\exp(\mathbf{f}_{i}\cdot\mathbf{f}_{k}/\tau)}\right),$ (5)
where $\mathbf{f}_{i},\mathbf{f}_{j}$ are features from the same domains. $y_{i}$ is the domain label of $\mathbf{f}_{i}$, and it is allocated with a unique value to each real/fake and previous/new domain. $\mathbb{I}_{[y_{i}\neq y_{k}]}$ denotes an indicator function, which equals 1 if $y_{i}\neq y_{k}$, and 0 otherwise. Notably, $\mathbf{f}$ could be the feature of current training data if they are from the new task, and generated by SUR or re-filled data if they are from the previous tasks. Meanwhile, to encourage the learning of diverse real domains, the real data from different tasks is also assigned with different unique $y_{i}$.
@@ 3.4.2 Incremental Decision Alignment
$\theta^{t+1}\leftarrow\left\|\theta^{t+1}\right\|_{2}\cdot\frac{(1-\gamma)\tilde{\theta}^{t+1}+\gamma\tilde{\theta}^{t}}{\left\|(1-\gamma)\tilde{\theta}^{t+1}+\gamma\tilde{\theta}^{t}\right\|_{2}},$ (6)
where $\tilde{\theta}=\frac{\theta}{\|\theta\|_{2}}$ and $\gamma$ denotes the learning rate. During training on the $(t+1)$-th task, the classifier $\mathcal{C}^{t+1}$ is optimized following Eq. 6 to be aligned with $\mathcal{C}^{t}$, while all previous classifiers are frozen to maintain the previous decision boundaries and their alignment.
@@ 3.5 Training and Inference
$y_{\text{infer}}=\sum_{i=1}^{t+1}\frac{\mathcal{C}^{i}(\mathbf{f})}{t+1}.$ (10)
Method Replays Task Protocol 1 Protocol 2
@@ 4.2 Comparisons with Existing Methods for Incremental Face Forgery Detection
In supplementary material, we provide results based on Protocol 3, which also indicates the superior performance of our method.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
We would like to thank all the reviewers for their constructive comments. Our work was supported by the National Natural Science Foundation of China (NSFC) under Grant No.62171324, No.62371350, and No.62372339.
@@
1 Further Results Comparing with SoTA
1.1 Results with Protocol 3
Method FF++ DFDCP DFD CDF Avg.
LwF [22] 67.34 67.43 84.05 87.90 76.68
CoReD [19] 74.08 76.59 93.41 80.78 81.22
DFIL [30] 86.28 79.53 92.36 83.81 85.49
DMP [41] 91.61 84.86 91.81 91.67 89.99
Ours 90.89 89.33 93.97 94.34 92.13
In Tab. 4, we copy the results after all tasks are incremented with P3 from their official papers [30, 41] to further compare the IFFD performance. Despite the notable distinction in experimental settings among these methods, our method still exhibits superior performance.
Method SDv21 FF++ DFDCP Avg.
Lower Bound 47.19 32.75 16.40 32.11
LwF 38.45 14.20 0.41 17.69
CoReD 12.80 11.21 2.05 8.69
DFIL 6.72 20.69 11.80 13.07
HDP 9.43 14.68 3.25 9.12
Ours 0.28 10.06 0.94 3.09
[…]
1.2 Evaluation with Forgetting Rate
Following [25], we compute FR based on AUC between current and first-learned models. Specifically, FR is calculated as $FR=1-\frac{AUC_{last}}{AUC_{first}}$, where $AUC_{last}$ is the AUC of one dataset tested on the currently-trained model, $AUC_{first}$ is the AUC of the model that firstly-introduced the dataset. The FR results in Tab. 5 indicate that our method has effectively tackled the issue of forgetting.
@@ 3.1 Generalization to Other Unseen Datasets
3.2 Generalization to Other Backbone
We additionally deployed our method on two mainstream backbones (ResNet and Xception) and compared the results with those of the original backbones under the same replay size. As shown in Tab. 7, our method also significantly improves the performance of these backbones.
Method SDv21 FF++ DFDCP CDF Avg.
Xception+Ours 0.996↑65.8% 0.767↑24.9% 0.852↑13.6% 0.951↑0.75% 0.892↑22.6%
ResNet+Ours 0.993↑85.8% 0.688↑16.0% 0.861↑20.0% 0.935↑0.73% 0.869↑25.4%
Table 7: Generalization to other backbones (AUC). $\uparrow$ denotes the improvement compared with vanilla backbones.